two capacitors of capacitance 1 and 2 respectively can be charged to a maximum voltage of 6kilovolt and 4 kilovolt.Find the maximum voltage that could be supplied to them whem they a re connected in series

Grade:12

1 Answers

Prakhar Shukla
6 years ago
Total charge on two capacitors=C1V1+C2V2=1×10^-6×6×10^3+8×10^-3=14×10^-3CIn series combination charge is conserved.Thus, charge on either capacitorsQ=14×10-3CV(1 microF)=q/C1=14kvV(2micro F)=q/C2=7kv